# DeepVigor+: Scalable and Accurate Semi-Analytical Fault Resilience Analysis for Deep Neural Networks

### v1

@@ III-E Stratified Random Sampling for Vulnerability Analysis:
The DeepVigor+ analysis can be conducted for all neurons/filters in a CNN. However, performing a complete analysis is obstructive and time-consuming, in particular for very deep and emerging CNNs. To address the scalability issue in resilience analysis for huge CNNs, we exploit the stratified sampling concept to reduce the sampling size for statistical DeepVigor+ analysis. Stratified random sampling is a method to increase the accuracy of estimation in sampling by dividing the population into subgroups called strata and random samples can be selected within them [48].
In DeepVigor+, the analysis is performed layer by layer. To take the benefits of stratified sampling for reducing the execution time of the analysis, we assume that each output channel in a layer is one stratum. Thereafter, a portion of all output channels specified by channel_sampling_ratio is considered and some random channels are selected. Within each channel, the number of analyzed neurons is determined by the $log_{2}(\#neurons)$ in the target channel, which is selected randomly. Therefore, the number of analyses can be described with Eq. (16).
#$\#analyzed\;neurons=\sum_{l\in layers}\lceil channel\_sampling\_ratio\times$
@@ IV Experimental Setup
IV-A DeepVigor+ Implementation
DeepVigor+ is fully implemented using Python and the Pytorch library. The source code of DeepVigor+ is fully open-source in https://github.com/mhahmadilivany/DeepVigor as a tool to enable researchers and engineers to adopt it for the resilience analysis of DNNs. The user can obtain the VF for the CONV layers of a target CNN based on analyzing neurons or weights by determining it through some inputs. User can specify the following inputs for the analysis:
  • Target pre-trained CNN: the tool loads the pre-trained set of weights.
  • Dataset: the tool loads the validation data from the dataset.
  • Analysis method: the tool performs resilience analysis for neurons (OFMaps) or filters (weights).
  • Sampling method: the tool performs either complete analysis or stratified random sampling based on channel sampling ratio.
  • Channel sampling ratio: in the case of stratified random sampling, the channel sampling ratio should be specified.
With the determined inputs, the tool performs the analysis and outputs the Layer Vulnerability Factor (LVF) for each layer and the Model Vulnerability Factor (MVF) for the entire CNN model. In the paper, we perform the DeepVigor+ analysis on one batch of 100 images in the test set. Note, that all experiments consider 32-bit floating-point data representation.
@@ IV-B Validating DeepVigor+
We use Fault Injection (FI) to validate the results of DeepVigor+. To that end, we validate the VF for channels by analyzing filters of CNNs using complete DeepVigor+ and weights FI. In this regard, before an inference, a random 3D filter in a specified layer is selected. In the FI campaign, one bit of one weight in the target filter is flipped considering 32-bit floating-point data representation and the inference is performed on the same data as the DeepVigor+ analysis was conducted. The same FI process is carried out for all bits of weights in the target filter resulting in the Channel Vulnerability Factor (CVF) calculated by the ratio of output misclassifications compared to the golden outcomes. The obtained CVF in FI for each channel is compared with the one in complete DeepVigor+ for the corresponding channel, and their absolute errors are reported. In all DNNs under the experiment, 15% of all channels throughout the CONV layers are passed through the FI campaign for validation.
To show the accuracy of the result in exploited stratified sampling in DeepVigor+, we compare the obtained VF for channels (CVF) and layers (LVF) in sampling DeepVigor+ vs. complete DeepVigor+ for both activations and filters. The absolute difference between obtained VFs is derived to show the VF estimation in sampling analysis. In order to achieve more accurate results, different channel_sampling_ratios are experimented including 5%, 10%, 15%, and 20%. Since neurons and channels are selected randomly in the stratified sampling, each sampled analysis is repeated 50 times and the maximum and mean absolute errors are reported.
@@
IV-C DNNs Under Study
In this paper, Deepvigor+ is executed and validated on six pre-trained DNNs using various datasets. DNNs under analysis include VGG-11 trained on CIFAR-10, VGG-16, ResNet-18-C and MobileNetV2 trained on CIFAR-100, ResNet-18-I and ResNet-34 trained on ImageNet. The baseline accuracy, number of channels and neurons for each DNN are shown in Table II. All experiments in this paper are performed on an NVIDIA A100 GPU accompanied by AMD EPYC 7742 64-core CPU.
@@ V-A DeepVigor+ Accuracy with FI
To analyze the accuracy of the obtained VFs by DeepVigor+, we perform fault injection experiments on 15% of randomly selected channels in all DNNs in Table II and present the Mean Absolute Error (MEA), as described in subsection IV-B. Table III presents the MEA results comparing the Channel Vulnerability Factor (CVF) by the complete DeepVigor+ filters analysis vs full FI into 15% of channels in each DNN. Noteworthy that the acceptable mean error is 1% [26].
The results in Table III indicate that the mean absolute error by DeepVigor+ compared to exact FI throughout the DNNs is between 0.819% to 0.978%, i.e., always less than 1%. These results demonstrate that DeepVigor+ is able to provide precise VFs and meets the expectation of an acceptable error with respect to exhaustive FI experiments.
Table III: Absolute error for CVF in DeepVigor+ and fault injection for 15% of the channels in DNNs.
DNN VGG-11 VGG-16
@@
The main sources of error in VF calculation in DeepVigor+ are:
@@
    As discussed in subsection III-C, DeepVigor+ approximates the error propagation in neurons based on the values of power of 2. This error approximation introduces an error to the resilience analysis which is also reflected in VF calculation.
@@
    DeepVigor+ assumes that bitflips resulting in big values are critical for DNNs and the distribution ratio for large values are always considered critical. However, in some cases, these values don’t result in misclassification. This phenomenon is another reason for a minor difference between the VF by FI and DeepVigor+.
@@
Table IV indicates the absolute error results over various channel sampling ratio for DNNs under study, in both neurons and filters analysis. Each sampling analysis is repeated 50 times to observe the effect of random selections in stratified sampling. Based on the results, the minimum absolute error throughout the experiments for both CVF and LVF is very close to 0. As observed, the difference between obtained CVFs and LVFs throughout the results is minimal, demonstrating the effectiveness of the exploited stratified sampling.
In all experiments, the Mean Absolute Error (MAE) for CVF does not vary since the sampling method within channels is similar (i.e., the logarithm of the number of OFMaps). Also, it is observed that the MEA CVF is less than 0.065% in all experiments for both activations and filters sampling analysis. It means that the averaged VF for the logarithm-based random sampling within each channel results in a highly accurate CVF. This phenomenon is a result of the unified distribution of weights within a channel in a pre-trained DNN. The maximum observed error in neuron analysis is 0.004% to 1.443% throughout DNNs, whereas the mean error remains low, meaning that for most of the channels, obtained CVFs are highly accurate leading to an overall high accuracy for obtained CVFs.
On the other hand, channel sampling ratio directly affects the accuracy of LVF calculations. As observed in Table IV, LVF error decreases with the increase of channel sampling ratio. Considering both MEA and maximum error for LVF, a 10% channel sampling ratio can guarantee a minimal error for VF calculations. Based on the MEA of complete DeepVigor+ compared to exhaustive FI Table III, sampling DeepVigor+ with $10\%$ channel sampling ratio ensures that all the overall error of obtained VFs for all DNNs will be lower than 1%.
In conclusion, the proposed stratified sampling in DeepVigor+ results in highly accurate VF for channels and layers obtained from both activations and filter analysis with a channel sampling ratio of 10%. Sampling DeepVigor+ results in a higher error than state-of-the-art statistic FI approaches such as data-aware and data-unaware, yet it meets the requirement of average error for resilience study which is 1%.
@@ V-C Run-Time and Scalability Investigation
To show the excellence of DeepVigor+ in terms of complexity, scalability and execution time against FI, first, we investigate the complexity of each based on the required number of simulations (i.e., forward pass executions) to obtain VF. Then we present the execution time for the complete and sampling DeepVigor+ on an NVIDIA A100 GPU.
Exhaustive FI is the most accurate method for determining precise VFs. In exhaustive FI, the required number of simulations equals the number of activations/weights times bit-width (i.e., 32 bits). Therefore, its complexity is proportional linearly to the size of DNNs. Whereas the complete DeepVigor+ analysis estimates VFs with high accuracy and significantly lower complexity. Although its complexity is affected by the size of DNNs, DeepVigor+ analysis exploits various optimizations to reduce the number of simulations resulting in significantly lower complexity than exhaustive FI. This is evidenced by the results in Table V where the required number of simulations is compared between Exhaustive FI and complete DeepVigor+ analysis, for activations and filters separately.
Table V: Number of simulations for exhaustive FI vs complete DeepVigor+ for activations and filters analysis.
@@
DNNs Exhaustive FI DeepVigor+ Exhaustive FI DeepVigor+
VGG-11 7,438,336 1,636,414 294,967,296 11,035
VGG-16 5,931,008 2,018,447 470,734,848 13,704
ResNet-18-C 21,331,968 3,387,189 357,095,424 15,906
MobileNet-V2 27,330,048 3,029,125 74,176,512 42,234
ResNet-18-I 69,844,992 20,730,314 357,341,184 23,450
ResNet-34 109,985,792 28,821,489 680,564,736 43,098
As observed, for each DNN, DeepVigor+ complete analysis requires significantly lower executions either in activations or filters analysis. Throughout the results, activations analysis by complete DeepVigor+ is 2.93 to 9.02 times faster than exhaustive FI. According to our detailed investigations, activations analysis by DeepVigor+ requires 6 forward simulations per neuron, on average, throughout the DNNs under study. In this regard, the introduced loss function to obtain the vulnerability values (box 4 in Fig 6) contributes to skipping the analysis for up to 26% of neurons. Moreover, the vulnerability value can be obtained by a single forward simulation for up to 55% of neurons due to separating VVSS exploration (boxes 7 and 8 in Fig. 6).
In the complete filters analysis, the DeepVigor+ fault propagation modeling implies a huge impact on the number of simulations in the orders of magnitude. DeepVigor+ can derive VF for filters 1,756 to 34,350 times faster than exhaustive FI. To obtain the vulnerability values for channels, up to 1% of channels are skipped by leveraging the loss function, and up to 34% of channels require one forward simulation. These results indicate that complete DeepVigor+ provides accurate VF for neurons and channels of CNNs with significantly lower complexity and shorter execution time than exhaustive FI enabled by its optimal fault propagation modeling and analysis.
Table VI: Comparison of required simulations for statistical FI [26] and sampling DeepVigor+.
@@
On the other hand, sampling DeepVigor+ is proposed to further reduce the execution time and complexity of resilience analysis and achieve a scalable method. To show its performance against statistical FI (SFI), Table VI compares the number of simulations for various state-of-the-art SFI [26] with sampling DeepVigor+. As presented, data-aware SFI leads to the least number of executions in FI-based simulation. It is observed that DeepVigor+ sampling activations analysis with $10\%$ channel sampling ratio leads to 8.72 to 20.5 times fewer simulations compared to data-aware SFI. For the filters analysis, DeepVigor+ obtains their VF with 59.4 up to 96.2 times fewer simulations than data-aware SFI.
To obtain the Model Vulnerability Factor (MVF) both activations and filters should be analyzed separately, based on Eq. (15). Therefore, sampling DeepVigor+ accelerates the process from 14.9 up to 26.9 times throughout the DNNs. The scalability and speed of the method are achieved by both channel sampling ratio and logarithmic sampling within them. It can be observed that with the remarkable growth of DNNs under analysis in their number of parameters, the number of simulations in DeepVigor+ does not grow linearly.
To demonstrate the execution time of DeepVigor+ analysis, we employed an A100 NVIDIA GPU and performed sampling DeepVigor+ for activations and filters with different channel sampling ratios. Table VII and Table VIII present the average execution time over 50 executions of the method for complete and sampling DeepVigor+ with different channel sampling ratios, for activations and filters respectively. It is observed that VFs can be obtained in a few minutes for any DNNs. Considering $10\%$ channel sampling ratio for both activations and filters analysis, the total MVF for VGG-11, VGG-16, ResNet-18-C, MobileNet-V2, ResNet-18-I and ResNet-34 is obtained almost in 8.7 minutes, 9.6 minutes, 12.7 minutes, 43.5 minutes, 19.4 minutes and 46 minutes, respectively. It is worth mentioning that the complete DeepVigor+ analysis for the DNNs under study takes almost 22 hours for VGG-16 and 18.5 days for ResNet-34, on the same GPU. Fast execution and accurate estimation of VF obtained by sampling DeepVigor+ analysis provide a remarkable opportunity for a high-speed resilience analysis for any DNN.
Table VII: Average execution time over 50 repetitions on A100 GPU for DeepVigor+ activations analysis, with different channel sampling ratios.
DNN 5% 10% 15% 20%
@@
Table VIII: Average execution time over 50 repetitions on A100 GPU for DeepVigor+ filters analysis, with different channel sampling ratios.
DNN 5% 10% 15% 20%
@@
V-D Reliability Visualization and Comparison for DNNs
It has been shown that VF for DNNs’ layers and the entire model can be obtained accurately in a few minutes by DeepVigor+. The obtained VF results by DeepVigor+ can be used to visualize the vulnerability of layers within a DNN and identify more vulnerable ones. Fig. 8 illustrates the LVF comparison for ResNet-18 trained on CIFAR-100 and ImageNet datasets, while total LVF for each layer is obtained based on the numerator of Eq. (15). This visualization sketches how much each layer is vulnerable to single faults compared to each other. As observed, in both ResNet18-C and ResNet18-I first layers are more vulnerable than the latter ones. Therefore, DeepVigor+ enables vulnerability visualization and comparison within a DNN.
Furthermore, DeepVigor+ results in total MVF based on Eq. (15) which is the weighted average of obtained LVFs providing a comprehensive examination of vulnerability between different DNNs. Fig. 9 indicates MVFs for activations and filters of DNNs, separately, as well as their total MVF. As a result, it is observed that activations are more vulnerable than weights. However, weights generally contribute more to the total MVF since their memory footprint is higher than that of activations in DNNs. Based on total MVF, VGG-16 is the least vulnerable DNN (MVF = 1.19%) and MobileNet-V2 is the most vulnerable one (MVF = 2.76%).
(a) ResNet-18-C
(b) ResNet-18-I
Figure 8: LVF visualization and comparison for ResNet-18 trained on a) CIFAR-100 and b) ImageNet.
VGG-11VGG-16ResNet-18-CMobileNetV2ResNet-18-IResNet-34001111222233334444Model Vulnerability Factor (%)MVF activationsMVF weightsMVF total
Figure 9: MVF comparison for DNNs based on activations, filters and the entire model derived by DeepVigor+.

### v2

@@ III-E Stratified Random Sampling for Vulnerability Analysis:
The DeepVigor+ analysis is capable of analyzing all neurons and filters in a CNN. However, performing a complete analysis is computationally complex and prohibitively time-consuming, in particular for very deep CNNs. Inspired by statistical FI, random sampling is the main solution for reducing the number of simulations, while preserving acceptable accuracy. However, naive random sampling of neurons across the entire CNN model or within individual layers can result in poor representativeness and accuracy of the resilience analysis, since the total population is extremely large.
To address the scalability issue in resilience analysis for huge CNNs without sacrificing accuracy, we exploit the concept of stratified random sampling to reduce the population size for statistical DeepVigor+ analysis. Stratified random sampling is a method to enhance the statistical estimation accuracy by dividing the population into meaningful subgroups or strata, and performing random sampling within them [50]. This method first randomly selects a subset of strata and then samples from each selected stratum to ensure balanced and representative coverage.
In DeepVigor+, the analysis is performed in a layer-wise manner, while each layer consists of multiple channels. We consider each output channel in a layer as a stratum, representing different features of the CNN. A portion of these output channels is randomly selected, specified by channel_sampling_ratio. Within each stratum (i.e., output channel), a subset of neurons is then randomly selected for analysis, determined by the logarithm of the total neuron count in the channel. This strategy ensures that the analysis remains computationally tractable while maintaining representative population coverage. The total number of analyses conducted under this approach can be defined in Eq. (16).
$\#analyzed\;neurons=\sum_{l\in layers}\lceil channel\_sampling\_ratio\times$
@@ IV Experimental Setup
IV-A DeepVigor+ Tool Implementation
DeepVigor+ is fully implemented using Python and the PyTorch library. The source code of DeepVigor+ is fully open-source in https://github.com/mhahmadilivany/DeepVigor as a tool to enable researchers and engineers to adopt it for the resilience analysis of DNNs. The user can obtain the VF for the CONV layers of a target CNN based on analyzing neurons or weights by determining it through some inputs. The user can specify the following inputs for the analysis:
  • model: The tool loads the pre-trained set of weights. By specifying the names of the model, it automatically loads the pretrained model. All the pretrained models from the open-source PyTorch CIFAR Models repository on GitHub [^1] https://github.com/chenyaofo/pytorch-cifar-models/tree/master and the PyTorch hub are integrated and supported within the tool.
  • dataset: The tool loads the test data from the dataset. It supports CIFAR-10, CIFAR-100, and ImageNet datasets.
  • batch-size: Specifies the batch-size of data to be analyzed by the tool.
  • run-method: The tool performs resilience analysis for neurons (OFMaps) or filters (weights). The user specifies the analysis method:
    • full-analysis-weight: Performs complete analysis on all weights of the CNN model,
    • full-analysis-act: Performs complete analysis on all activations of the CNN model,
    • sampling-analysis-weight: Performs sampling analysis on the logarithm number of weights per randomly selected channel based on channel sampling ratio,
    • sampling-analysis-act: Performs sampling analysis on the logarithm number of activations per randomly selected channel based on channel sampling ratio,
  • Channel-sampling-ratio: In the case of sampling analysis, this parameter specifies the percentage of the random channels in each CONV layer to be analyzed.
With the determined inputs, the tool performs the analysis based on the presented method in Section III and outputs the Layer Vulnerability Factor (LVF) for each layer. The tool creates an individual directory for a pretrained model. In the directory, the tool creates a log file and names it based on the specified model, dataset, run-mode, and channel-sampling-ratio, and saves the obtained VF results. Therefore, the Model Vulnerability Factor (MVF) for the entire CNN model can be calculated based on VF for activations and weights. In the paper, we perform the DeepVigor+ analysis on one batch of 100 images in the test set. Note that all experiments consider 32-bit floating-point data representation.
@@ IV-B Validating DeepVigor+
We use Fault Injection (FI) to validate the results of DeepVigor+. To that end, we validate the VF for channels by analyzing filters of CNNs using complete DeepVigor+ and weights FI. In this regard, before an inference, a random 3D filter in a specified layer is selected. In the FI campaign, one bit of one weight in the target filter is flipped, considering 32-bit floating-point data representation, and the inference is performed on the same data as the DeepVigor+ analysis was conducted. The same FI process is carried out for all bits of weights in the target filter, resulting in the Channel Vulnerability Factor ($CVF_{FI}$) calculated by the ratio of changes in the golden classifications. $CVF_{FI}$ for each channel is compared with the $CVF_{weight}$ for the same channel in complete DeepVigor+, and their absolute errors are reported. In all DNNs under the experiment, $15\%$ of all channels throughout the CONV layers are passed through the FI campaign for validation.
To show the accuracy of the results in stratified sampling in DeepVigor+, we compare the obtained VF for channels (CVF) and layers (LVF) in sampling DeepVigor+ vs. complete DeepVigor+ for activations and filters separately. The absolute difference between the obtained VFs is derived to show the VF estimation in sampling analysis. Different channel_sampling_ratios are explored including $5\%$, $10\%$, $15\%$, and $20\%$. Since neurons and channels are selected randomly in the stratified sampling, each sampling analysis is repeated $50$ times, and their maximum and Mean Absolute Error (MAE) are reported.
@@
IV-C CNNs Under Study
In this paper, DeepVigor+ is executed and validated on six pre-trained DNNs using various datasets. CNNs under analysis include VGG-11 trained on CIFAR-10, VGG-16, ResNet-18-C, and MobileNetV2 trained on CIFAR-100, ResNet-18-I and ResNet-34 trained on ImageNet. The baseline accuracy, number of channels and neurons for each CNN are shown in Table II. All experiments in this paper are performed on an NVIDIA A100 GPU accompanied by an AMD EPYC 7742 64-core CPU.
@@ V-A DeepVigor+ Accuracy with FI
To analyze the accuracy of the obtained VFs by DeepVigor+, we perform exhaustive FI experiments on $15\%$ of randomly selected channels in all CNNs in Table II and compare the results with complete DeepVigor+, as described in subsection IV-B. Table III presents the MAE between $CVF_{FI}$ and $CVF_{weight}$ by the complete DeepVigor+’s filters analysis for each CNN. The acceptable mean error for a reliability analysis compared to exhaustive FI is $1\%$ [27].
The results in Table III indicate that the MAE by DeepVigor+ compared to exact FI throughout the CNNs is between $0.819\%$ to $0.978\%$, i.e., always less than $1\%$. These results demonstrate that DeepVigor+ is able to provide precise VFs and meets the expectation of an acceptable error with respect to exhaustive FI experiments.
Table III: MAE of VF for weight channels between complete DeepVigor+ and exhaustive FI for $15\%$ of the channels in CNNs.
CNN VGG-11 VGG-16
@@
The main sources of error in DeepVigor+ are as follows:
@@
    As discussed in subsection III-C, DeepVigor+ approximates the error propagation in neurons based on the values of power of $2$. This error approximation introduces an error to the resilience analysis, which is also reflected in the VF calculation.
@@
    DeepVigor+ assumes that bitflips resulting in big values are critical for CNNs, and the distribution ratio for large values is always considered critical. However, in some cases, these values don’t result in the change of the golden classification. This phenomenon is another reason for a minor difference between the VF by FI and DeepVigor+.
@@
Table IV indicates the absolute error results over various channel sampling ratios for CNNs under study, in both neurons and filters analysis. Each sampling analysis is repeated $50$ times to observe the effect of random selections in stratified sampling. Based on the results, the minimum absolute error throughout the experiments for both CVF and LVF is very close to 0. As observed, the difference between obtained CVFs and LVFs throughout the results is minimal, demonstrating the effectiveness of the exploited stratified sampling.
In all experiments, the MAE for CVF does not vary since the sampling method within channels is similar (i.e., the logarithm of the number of OFMaps). Also, it is observed that the MAE CVF is less than $0.065\%$ throughout the experiments for both activations and filters sampling analysis. It means that the averaged VF for the logarithm-based random sampling within each channel results in a highly accurate CVF. This phenomenon is a result of the unified distribution of weights within a channel in a pre-trained CNN. The maximum observed error in neuron analysis is $0.004\%$ to $1.443\%$ throughout CNNs, whereas the mean error remains low, meaning that for most of the channels, obtained CVFs are highly accurate, leading to an overall high accuracy for CVFs.
On the other hand, channel sampling ratio directly affects the accuracy of LVF calculations. As observed in Table IV, LVF error decreases with the increase of channel sampling ratio. Considering both MAE and maximum error for LVF, a $10\%$ channel sampling ratio can guarantee a minimal error for VF calculations. Based on the MAE of complete DeepVigor+ compared to exhaustive FI Table III, sampling DeepVigor+ with a $10\%$ channel sampling ratio ensures that all the overall errors of the obtained VFs for all CNNs will be lower than $1\%$.
In conclusion, the proposed stratified sampling in DeepVigor+ results in highly accurate VF for channels and layers obtained from both activations and filter analysis with a channel sampling ratio of $10\%$. Sampling DeepVigor+ results in a higher error than state-of-the-art statistical FI approaches, such as data-aware and data-unaware, yet it meets the above-mentioned requirement of average error for the resilience study, which is $1\%$.
@@ V-C Run-Time and Scalability Investigation
To show the effectiveness of DeepVigor+ in terms of complexity, scalability, and execution time against DeepVigor and FI, first, we investigate the complexity of each method based on the required number of simulations (i.e., forward pass executions) to obtain VF. Then we present the execution time for the complete and sampling DeepVigor+ on an NVIDIA A100 GPU.
Exhaustive FI is the most accurate method for determining precise VFs. In exhaustive FI, the required number of simulations equals the number of activations/weights times bit-width (i.e., 32 bits). Therefore, its complexity is proportional linearly to the size of CNNs. On the other hand, DeepVigor [8] is faster than exhaustive FI to obtain the VF for activations. To obtain a vulnerability value range, DeepVigor requires $9.1$ simulations per neuron, on average.
Whereas the complete DeepVigor+ analysis estimates VFs with high accuracy and significantly lower complexity. Although its complexity is affected by the size of CNNs, DeepVigor+ analysis exploits various optimizations to reduce the number of simulations, resulting in significantly lower complexity than exhaustive FI and DeepVigor. This is evidenced by the results in Table V where the required number of simulations is compared between Exhaustive FI, DeepVigor, and complete DeepVigor+ analysis, for activations and filters separately.
Table V: Number of simulations for exhaustive FI, DeepVigor, and complete DeepVigor+ for activations and filters analysis.
@@
CNNs Exhaustive FI DeepVigor Complete DeepVigor+ Exhaustive FI Complete DeepVigor+
VGG-11 7,438,336 5,124,816 1,636,414 294,967,296 11,035
VGG-16 5,931,008 3,586,312 2,018,447 470,734,848 13,704
ResNet-18-C 21,331,968 9,941,732 3,387,189 357,095,424 15,906
MobileNet-V2 27,330,048 12,958,887 3,029,125 74,176,512 42,234
ResNet-18-I 69,844,992 45,083,325 20,730,314 357,341,184 23,450
ResNet-34 109,985,792 66,303,692 28,821,489 680,564,736 43,098
As observed, for each CNN, DeepVigor+ complete analysis requires significantly fewer simulations than exhaustive FI and DeepVigor, in both activations and filters analysis. Throughout the results, activation analysis by complete DeepVigor+ is $2.93$ to $9.02$ times faster than exhaustive FI. Furthermore, DeepVigor+ outperforms DeepVigor in activations analysis, requiring $1.77$ to $4.28$ times fewer simulations. Note that the accuracy of DeepVigor to obtain VF is $0.1\%$ while the accuracy of DeepVigor+ is $1\%$.
According to our detailed investigations, activation analysis by DeepVigor+ requires $6$ forward simulations per neuron, on average, throughout the CNNs under study. In this regard, the introduced loss function to obtain the vulnerability values (box 4 in Fig. 6) contributes to skipping the analysis for up to $26\%$ of neurons. Moreover, the vulnerability value can be obtained by a single forward simulation for up to $55\%$ of neurons due to separating VVSS exploration (boxes 7 and 8 in Fig. 6).
Moreover, in the complete filters analysis, the DeepVigor+ fault propagation modeling implies an even more significant impact on the number of simulations in the orders of magnitude. DeepVigor+ can derive VF for filters $1,756$ to $34,350$ times faster than exhaustive FI. To obtain the vulnerability values for channels, up to $1\%$ of channels are skipped by leveraging the loss function, and up to $34\%$ of channels require one forward simulation. These results indicate that complete DeepVigor+ provides accurate VF for neurons and channels of CNNs with significantly lower complexity and shorter execution time than exhaustive FI enabled by its optimal fault propagation modeling and analysis.
Table VI: Comparison of required simulations for different statistical FI methods based on [27] and sampling DeepVigor+.
@@
On the other hand, sampling DeepVigor+ is proposed to further reduce the execution time and complexity of resilience analysis and achieve a scalable method. To show its performance against Statistical FI (SFI), Table VI compares the number of simulations for various state-of-the-art SFI methods proposed in [27] with sampling DeepVigor+. As presented, data-aware SFI leads to the least number of executions in the FI-based simulation. It is observed that DeepVigor+ sampling activations analysis with $10\%$ channel sampling ratio leads to $8.72$ to $20.5$ times fewer simulations compared to data-aware SFI. For the filters analysis, DeepVigor+ obtains their VF with $59.4$ up to $96.2$ times fewer simulations than data-aware SFI, which is currently accepted to be the fastest state-of-the-art approach.
To obtain the Model Vulnerability Factor (MVF) both activations and filters should be analyzed separately, based on Eq. (15). Therefore, sampling DeepVigor+ accelerates the process, ranging from $14.9$ up to $26.9$ times throughout the CNNs. The scalability and speed of the method are achieved by both channel sampling ratio and logarithmic sampling within them. It can be observed that with the remarkable growth of CNNs under analysis in their number of parameters, the number of simulations in DeepVigor+ does not grow linearly.
To demonstrate the execution time of DeepVigor+ analysis, we employed an A100 NVIDIA GPU and performed sampling DeepVigor+ for activations and filters with different channel sampling ratios. Table VII and Table VIII present the average execution time over $50$ executions of the method for complete and sampling DeepVigor+ with different channel sampling ratios, for activations and filters, respectively. It is observed that VFs can be obtained in a few minutes for any CNN.
Considering $10\%$ channel sampling ratio for both activations and filters analysis, the total MVF for VGG-11, VGG-16, ResNet-18-C, MobileNet-V2, ResNet-18-I and ResNet-34 is obtained in almost $8.7$ minutes, $9.6$ minutes, $12.7$ minutes, $43.5$ minutes, $19.4$ minutes, and $46$ minutes, respectively. It is worth mentioning that the complete DeepVigor+ analysis for the CNNs under study takes almost $22$ hours for VGG-16 and $18.5$ days for ResNet-34, on the same GPU. Fast execution and accurate estimation of VF obtained by sampling DeepVigor+ analysis provide a remarkable opportunity for a high-speed resilience analysis for any known CNN.
Table VII: Average execution time on A100 GPU for DeepVigor+ activations analysis, with different Channel Sampling Ratios (CSR).
CNN CSR = 5% CSR = 10% CSR = 15% CSR = 20%
Neurons
@@
Table VIII: Average execution time on A100 GPU for DeepVigor+ filters analysis, with different Channel Sampling Ratios (CSR).
CNN CSR = 5% CSR = 10% CSR = 15% CSR = 20%
Filters
@@
Furthermore, the implementation of DeepVigor+ is resource-efficient and does not impose significant GPU memory demands. Across all experiments, using a $10\%$ channel sampling ratio and a batch-size of $100$, the GPU memory consumption ranges from a minimum of $4$ GB to a maximum of $5$ GB. This demonstrates both the computational efficiency and scalability of the proposed method. As a result, DeepVigor+ can be readily deployed on a wide range of hardware platforms, including GPUs with limited memory capacity, making it accessible for broader adoption in both academic and industrial settings.
V-D Vulnerability Visualization and Comparison for CNNs
It has been demonstrated that VF for CNN layers and the entire model can be obtained accurately in a few minutes by DeepVigor+. The obtained VF results by DeepVigor+ can be used to visualize the vulnerability of layers within a CNN and identify the most vulnerable ones, leading to efficient protection mechanisms. Fig. 8 illustrates the comparison of total LVF for ResNet-18 trained on ImageNet, while total LVF for each layer is obtained based on the numerator of Eq. (15), which considers the LVF of both weights and activations. This visualization sketches how much each layer is vulnerable to single faults compared to the others. As observed, the first layers are generally more vulnerable than the latter ones. Layers 1 to 6, 8, and 13 demonstrate more vulnerability, and they need stronger fault mitigation techniques, while in other layers, the costs of implementing fault tolerance can be saved.
Figure 8: LVF visualization and comparison for ResNet-18 on ImageNet.
Furthermore, DeepVigor+ results in total MVF based on Eq. (15), which is the weighted average of obtained LVFs, providing a comprehensive examination of vulnerability between different CNNs. Fig. 9 indicates MVFs for activations and filters of CNNs, separately, as well as their total MVF. As a result, it is observed that activations are more vulnerable than weights. However, weights generally contribute more to the total MVF since their number is higher than that of activations in CNNs. Except in MobilNetV2, the number of weights is smaller, so the activations have a more significant impact on the total MVF. Based on total MVF, VGG-16 is the least vulnerable CNN (MVF = $1.19\%$), and MobileNetV2 is the most vulnerable one (MVF = $2.76\%$).
This visualization also enables analyzing the impact of architectural characteristics of CNNs on their fault resilience. As observed, in both VGG and ResNet, their deeper versions, VGG-16 and ResNet-34, possess a smaller MVF than their shallower counterparts, VGG-11 and ResNet-18-I, respectively. Also, the specific architecture of MobileNet-V2, which contains inverted residual blocks and depthwise separable convolutions, leads to more vulnerability than other CNN architectures.
VGG-11VGG-16ResNet-18-CMobileNetV2ResNet-18-IResNet-34011223344Model Vulnerability Factor (%)MVF activationsMVF weightsMVF total
Figure 9: MVF comparison for CNNs based on activations, filters, and the entire model derived by DeepVigor+.